# RI-MAE: Rotation-Invariant Masked AutoEncoders for Self-Supervised Point Cloud Representation Learning

### v1

@@ Abstract
Masked point modeling methods have recently achieved great success in self-supervised learning for point cloud data. However, these methods are sensitive to rotations and often exhibit sharp performance drops when encountering rotational variations. In this paper, we propose a novel Rotation-Invariant Masked AutoEncoders (RI-MAE) to address two major challenges: 1) achieving rotation-invariant latent representations, and 2) facilitating self-supervised reconstruction in a rotation-invariant manner. For the first challenge, we introduce RI-Transformer, which features disentangled geometry content, rotation-invariant relative orientation and position embedding mechanisms for constructing rotation-invariant point cloud latent space. For the second challenge, a novel dual-branch student-teacher architecture is devised. It enables the self-supervised learning via the reconstruction of masked patches within the learned rotation-invariant latent space. Each branch is based on an RI-Transformer, and they are connected with an additional RI-Transformer predictor. The teacher encodes all point patches, while the student solely encodes unmasked ones. Finally, the predictor predicts the latent features of the masked patches using the output latent embeddings from the student, supervised by the outputs from the teacher. Extensive experiments demonstrate that our method is robust to rotations, achieving the state-of-the-art performance on various downstream tasks.
Introduction
@@
Pioneered by PointNet (Qi et al. 2017a), deep neural networks have achieved great success in point cloud representation learning, and shown impressive results on various downstream tasks such as point cloud classification and segmentation (Qi et al. 2017b; Wang et al. 2019; Li et al. 2018). However, such approaches are supervised and necessitate data annotated with labor-intensive efforts for training models. In contrast, self-supervised learning methods learn to derive representation from unlabeled data by designing pretext tasks, achieving notable success in the fields of natural language processing (NLP) and computer vision (CV). Inspired by this success, many self-supervised point cloud representation learning approaches have been proposed (Yang et al. 2018; Liu et al. 2019; Afham et al. 2022; Chen et al. 2021; Du et al. 2021; Poursaeed et al. 2020; Rao, Lu, and Zhou 2020; Sanghi 2020; Tsai et al. 2022; Sun et al. 2022). They learn to acquire point cloud representations through a variety of tasks, such as diverse forms of point cloud reconstruction and contrastive learning tasks.
@@
    How to achieve rotation-invariant latent representations? Existing methods attain rotation invariance by transforming the input point cloud into handcrafted features (Chen et al. 2019; Sun, Lian, and Xiao 2019; Li et al. 2021b), designing rotation-invariant convolutions (Chen and Cong 2022; Zhang, Hua, and Yeung 2022), or establishing global reference frames (Zhao et al. 2022; Li et al. 2021a). Nevertheless, directly integrating these solutions into transformer-based MPM poses significant challenges, as they demand specialized network designs that are inherently incompatible with the transformer framework. Despite the introduction of a rotation-invariant transformer in (Yu, Zhang, and Cai 2023), its complex architecture hinders the development of a universally applicable MPM framework. Hence, instead of making extensive modifications to the Transformer, we aim to achieve rotation invariance with minimal modifications, preserving the advantages of transformer-based MPM architectures.
@@
Related Work
Rotation-Invariant Point Cloud Analysis
@@
To learn consistent features across a range of potential rotations, many methods (Chen et al. 2019; Sun, Lian, and Xiao 2019; Li et al. 2021b; Gu et al. 2021) transform the input point cloud into rotation-invariant, handcrafted features, such as distances and angles. These features, derived from point coordinates and normal vectors, remain independent regarding the global orientation of point clouds. Drawing inspiration from 2D image representation learning, several methods (Chen and Cong 2022; Zhang, Hua, and Yeung 2022; Zhang et al. 2020a) introduced rotation-invariant convolutions to extract point cloud features. However, by solely focusing on distances and angles, these methods may lose other vital information. To address this, some methods align point clouds to canonical poses, thereby learning rotation-invariant representations through the establishment of global reference frames (Li et al. 2021a; Zhang et al. 2020b; Lou et al. 2023; Liu et al. 2022). Unlike these existing approaches generally requiring complex feature engineering and network designs, our method eliminates the need for intricate feature or network designs with a simplified process, and allows for training without the necessity of annotating data.
@@
Figure 2: Overview of the proposed RI-MAE architecture. The input point cloud is first divided into point patches via FPS and KNN, and PCA is utilized to align the point patches and obtain rotation matrices relative to canonical poses. Then geometry content tokens, RI-OEs, and RI-PEs are formulated as RI-Transformer’s inputs. Finally, task heads are employed for downstream tasks. A dual-branch student-teacher scheme is devised to conduct the self-supervised learning to pretrain the RI-Transformer.
Self-Supervised Learning for Point Clouds
@@
Self-supervised learning for point clouds has been explored extensively, leading to many promising approaches. These approaches train models with carefully designed pretext tasks instead of using labeled data. Common pretext tasks include point cloud reconstruction efforts such as input reconstruction (Yang et al. 2018), local-to-global reconstruction (Liu et al. 2019), restoration of anomalous parts (Chen et al. 2021), and completion of occlusions (Wang et al. 2020). Additionally, various contrastive learning approaches have been explored (Du et al. 2021; Xie et al. 2020; Rao, Lu, and Zhou 2020; Sanghi 2020; Afham et al. 2022), which learn point cloud representations by differentiating between pre-defined positive and negative samples. Other pretext tasks for self-supervised learning include orientation estimation (Poursaeed et al. 2020), mixing and disentangling techniques (Sun et al. 2022), and pose disentanglement strategies (Tsai et al. 2022).
@@
Recently, inspired by the great achievement of BERT (Devlin et al. 2018) and MAE (He et al. 2021), many Transformer-based MPM methods for point cloud have been proposed and significantly outperform other methods in downstream tasks. Point-BERT (Yu et al. 2022) introduced the BERT pretraining scheme to point clouds, which trains models with a task to predict discrete tokens generated by dVAE (Rolfe 2016). Point-MAE (Pang et al. 2022) proposed a self-supervised pretext task to reconstruct masked input point patches based on the known tokens. Point-M2AE (Zhang et al. 2020a) employed a pyramid encoder and decoder design to conduct masked point cloud reconstruction in a hierarchical manner. ACT (Dong et al. 2023), I2P-MAE (Zhang et al. 2023), ReCon (Qi et al. 2023) and ReCon++ (Qi et al. 2024) improved the 3D masked modeling by introducing pre-trained model from other modalities, such as images and natural languages. Point-GPT (Chen et al. 2024a) arranged point patches in an ordered sqeuence and extended the concept of genreative pre-training transformer (GPT) to point clouds. Despite achieving remarkable results, these MPM methods are sensitive to rotation, constraining their scalability and generalizability.
Method
Revisiting Masked Point Modeling
@@
We first revisit the masked point modeling (MPM) framework for point clouds. Existing methods, such as Point-MAE and Point-M2AE, segment the input point cloud $\mathbf{X}$ into $G$ point patches $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{G}$ using Farthest Point Sampling (FPS) and K-Nearest Neighborhood (KNN) algorithms. Subsequently, a masking ratio $\alpha$ is applied, partitioning the point patches into a visible set $\mathbf{P}_{v}=\{\mathbf{p}_{i}^{v}\}_{i=1}^{V}$ and a masked set $\mathbf{P}_{m}=\{\mathbf{p}_{i}^{m}\}_{i=1}^{M}$, where $V=(1-\alpha)\times G$ and $M=\alpha\times G$ are the numbers of visible and masked patches, respectively. Following this, a mini-PointNet (Qi et al. 2017a) is utilized to tokenize the visible point patches as point tokens $\mathbf{T}_{v}=\{\mathbf{t}_{i}^{v}\}_{i=1}^{V}$ and the coordinates of patch centers are projected as position embeddings $\{\mathbf{pos}_{i}\}_{i=1}^{G}$ via a learnable MLP (multi-layer perceptron). To this end, an encoding stage is introduced, during which only the visible tokens $\mathbf{T}_{v}$ are fed into a transformer-based encoder:
@@
Rotation-Invariant Transformer
@@
$\bar{\mathbf{p}}_{i}^{c},\mathbf{R}_{i}^{\prime}=\mathrm{PCA}(\mathbf{p}_{i}\mathbf{R}),\mathbf{R}_{i}^{\prime}=\mathbf{R}_{i}\mathbf{R},$ (7)
where $\mathbf{R}_{i}^{\prime}$ is the rotation that aligns the rotated point patch $\mathbf{p}_{i}\mathbf{R}$ into the canonical $\bar{\mathbf{p}}_{i}$. Hence, the orientation embedding $\mathbf{ori}_{i}$ is affected as well, since the MLP itself does not inherently ensure rotation invariance.
@@
where $\mathbf{Q}=\bar{\mathbf{T}}\mathbf{W}^{Q}$, $\mathbf{K}=\bar{\mathbf{T}}\mathbf{W}^{K}$, $\mathbf{V}=\bar{\mathbf{T}}\mathbf{W}^{V}$ are the query, key and value matrices; $\mathbf{W}^{Q},\mathbf{W}^{K}$ and $\mathbf{W}^{V}$ are learnable matrices; and $\mathbf{B}=\{b_{ij}\}\in\mathbb{R}^{G\times G}$ is the relative orientation embedding. We formulate $b_{ij}$ by establishing bilinear relations that integrate the features of the $i^{\text{th}}$ patch token with the relative orientation information pertaining to the $j^{\text{th}}$ patch:
@@
Note that for computational efficiency, we adopt the query weight matrix for the bilinear operator empirically.
@@
where $\mathbf{c}_{i}\mathbf{R}_{i}^{T}$ represents the location of the origin in the local reference frame of $\mathbf{p}_{i}$.
@@
Rotation-Invariant MAE
Existing MPM methods utilize raw point clouds to guide the reconstruction of masked patches. However, this approach is not suitable for building a rotation-invariant MPM. Specifically, for a given point cloud in different directions, a rotation-invariant MPM should yield identical point cloud latent representations; however, the coordinates of the masked patches are different, necessitating that the decoder reconstructs distinct patches from the same latent representation. This conflict results in unstable loss, making the model difficult for optimization.
@@
Experiments
Implementation Details
@@
For training details, we utilized an AdamW optimizer with a cosine learning rate decay, applying a decay factor 0.05, and incorporated a 10-epoch warm-up phase. The initial learning rate was set at 0.0005 for both pretraining and classification tasks, while for segmentation tasks, it was adjusted to 0.0002. We pretrained our RI-MAE on the train set of ShapeNet (Chang et al. 2015), which consists of more than 50,000 3D models, covering 55 categories. Following ClusterNet (Chen et al. 2019), we conducted experiments across three scenarios to evaluate the rotation invariance: 1) z/z for training and testing with azimuthal rotations; 2) z/SO3 for training with azimuthal rotations and testing with arbitrary rotations; and 3) SO3/SO3, entailing both training and testing under arbitrary rotations. All experiments were conducted utilizing two RTX 2080Ti GPUs, with the PyTorch framework version 1.7.
Downstream Tasks
@@
We evaluate the proposed method on a classification task with a real-word dataset - ScanOjectNN (Uy et al. 2019), which contains 2,902 point clouds of 15 categories collected in real world. There are three variants of ScanOjectNN: OBJ_BG, OBJ_ONLY, and PB_T50_RS. We evaluate our approach on the OBJ_BG variant by incorporating additional MLP classification heads, with 1,024 points sampled. The comparative results of our method and existing approaches are detailed in Table 1. Specifically, we trained an RI-Transfomer without utilizing the proposed dual-branch self-supervised learning strategy as a baseline. The baseline RI-Transformer model exhibits superior performance compared to the state-of-the-art rotation-invariant methods, notably outperforming the second-best approach (Yu, Zhang, and Cai 2023), which is also based on a transformer architecture, by a margin of +1.7%. Regrading RI-MAE, with the dual-branch learning strategy, it delivers improvements of +2.6% over the RI-Transfomer baseline. These results indicate the superiority of the proposed mechanisms.
@@
ModelNet40 is a synthetic shape dataset that contains 12,311 CAD models across 40 categories, with 1,024 points sampled from each model (Wu et al. 2014). We conducted few-shot learning on ModelNet40, following practices from existing MPM methods. The performance is assessed across two settings: 10-way 10-shot and 10-way 20-shot, where $K$-way $N$-shot indicates that $N$ objects are sampled for training from each of $K$ randomly selected classes, followed by testing on 20 unseen instances randomly chosen from each of these $K$ classes. All experiments were conducted in the z/SO3 scenario to evaluate rotation robustness and generalization capabilities. We report both the average accuracy and the standard deviation from 10 independent runs in Table 2. The results show that our RI-MAE significantly outperforms existing methods, showing more than 50% improvements.
@@
The evaluation on part segmentation is with ShapeNet, which consists of 16,881 objects across 16 categories. In line with existing MPM methods, we sample 2,048 points from each object as input. A simple segmentation head is incorporated as in Point-MAE (Pang et al. 2022). The segmentation results in the z/SO3 scenario are reported in Table 3. It is evident that existing MPM methods are sensitive to rotations of the point cloud, resulting in significantly unsatisfactory accuracy levels. In contrast, our RI-MAE demonstrates robustness to rotations and achieves the best performance, with a mean intersection over union (mIoU) score of 84.3% $\mathrm{mIoU}_{I}$ for instances and 82.1% $\mathrm{mIoU}_{C}$ for classes. Specifically, the baseline RI-Transfomer outperforms the state-of-the-art rotation-invariant methods as well and our RI-MAE further enhanced the capability. The part segmentation results of RI-MAE in the z/SO3 scenario are visualized in Fig. 3, indicating consistent predictions regardless of rotation.
@@
Table 5: Ablation study on the proposed RI-OE, RI-PE and dual-branch.

### v2

@@ Abstract
Masked point modeling methods have recently achieved great success in self-supervised learning for point cloud data. However, these methods are sensitive to rotations and often exhibit sharp performance drops when encountering rotational variations. In this paper, we propose a novel Rotation-Invariant Masked AutoEncoders (RI-MAE) to address two major challenges: 1) achieving rotation-invariant latent representations, and 2) facilitating self-supervised reconstruction in a rotation-invariant manner. For the first challenge, we introduce RI-Transformer, which features disentangled geometry content, rotation-invariant relative orientation and position embedding mechanisms for constructing rotation-invariant point cloud latent space. For the second challenge, a novel dual-branch student-teacher architecture is devised. It enables the self-supervised learning via the reconstruction of masked patches within the learned rotation-invariant latent space. Each branch is based on an RI-Transformer, and they are connected with an additional RI-Transformer predictor. The teacher encodes all point patches, while the student solely encodes unmasked ones. Finally, the predictor predicts the latent features of the masked patches using the output latent embeddings from the student, supervised by the outputs from the teacher. Extensive experiments demonstrate that our method is robust to rotations, achieving the state-of-the-art performance on various downstream tasks. Our code is available at https://github.com/kunmingsu07/RI-MAE .
1 Introduction
@@
Pioneered by PointNet (Qi et al. 2017a), deep neural networks have achieved great success in point cloud representation learning, and shown impressive results on various downstream tasks such as point cloud classification and segmentation (Qi et al. 2017b; Wang et al. 2019; Li et al. 2018). However, such approaches are supervised and necessitate data annotated with labor-intensive efforts for training models. In contrast, self-supervised learning methods learn to derive representation from unlabeled data by designing pretext tasks, achieving notable success in the fields of natural language processing (NLP) and computer vision (CV). Inspired by this success, many self-supervised point cloud representation learning approaches have been proposed (Yang et al. 2018; Liu et al. 2019; Afham et al. 2022; Chen et al. 2021; Poursaeed et al. 2020; Rao, Lu, and Zhou 2020; Sanghi 2020; Tsai et al. 2022; Sun et al. 2022). They learn to acquire point cloud representations through a variety of tasks, such as diverse forms of point cloud reconstruction and contrastive learning tasks.
@@
    How to achieve rotation-invariant latent representations? Existing methods attain rotation invariance by transforming the input point cloud into handcrafted features (Chen et al. 2019; Li et al. 2021b), designing rotation-invariant convolutions (Chen and Cong 2022; Zhang, Hua, and Yeung 2022), or establishing global reference frames (Zhao et al. 2022; Li et al. 2021a). Nevertheless, directly integrating these solutions into transformer-based MPM poses significant challenges, as they demand specialized network designs that are inherently incompatible with the transformer framework. Despite the introduction of a rotation-invariant transformer in (Yu, Zhang, and Cai 2023), its complex architecture hinders the development of a universally applicable MPM framework. Hence, instead of making extensive modifications to the Transformer, we aim to achieve rotation invariance with minimal modifications, preserving the advantages of transformer-based MPM architectures.
@@
2 Related Work
2.1 Rotation-Invariant Point Cloud Analysis
@@
To learn consistent features across a range of potential rotations, many methods (Chen et al. 2019; Li et al. 2021b; Gu et al. 2021) transform the input point cloud into rotation-invariant, handcrafted features, such as distances and angles. These features, derived from point coordinates and normal vectors, remain independent regarding the global orientation of point clouds. Drawing inspiration from 2D image representation learning, several methods (Chen and Cong 2022; Zhang, Hua, and Yeung 2022) introduced rotation-invariant convolutions to extract point cloud features. However, by solely focusing on distances and angles, these methods may lose other vital information. To address this, some methods align point clouds to canonical poses, thereby learning rotation-invariant representations through the establishment of global reference frames (Li et al. 2021a; Zhang et al. 2020; Lou et al. 2023). Unlike these existing approaches generally requiring complex feature engineering and network designs, our method eliminates the need for intricate feature or network designs with a simplified process, and allows for training without the necessity of annotating data.
@@
Figure 2: Overview of the proposed RI-MAE architecture. The input point cloud is first divided into point patches via FPS and KNN, and PCA is utilized to align the patches and obtain rotation matrices relative to canonical poses. Then geometry content tokens, RI-OEs, and RI-PEs are formulated as RI-Transformer’s inputs. Finally, task heads are employed for downstream tasks. A dual-branch student-teacher scheme is devised to conduct the self-supervised learning to pretrain the RI-Transformer.
2.2 Self-Supervised Learning for Point Clouds
@@
Self-supervised learning for point clouds has been explored extensively, leading to many promising approaches. These approaches train models with carefully designed pretext tasks instead of using labeled data. Common pretext tasks include point cloud reconstruction efforts such as input reconstruction (Yang et al. 2018), local-to-global reconstruction (Liu et al. 2019), restoration of anomalous parts (Chen et al. 2021), and completion of occlusions (Wang et al. 2020). Additionally, various contrastive learning approaches have been explored (Xie et al. 2020; Rao, Lu, and Zhou 2020; Sanghi 2020; Afham et al. 2022), which learn point cloud representations by differentiating between pre-defined positive and negative samples. Other pretext tasks for self-supervised learning include orientation estimation (Poursaeed et al. 2020), mixing and disentangling techniques (Sun et al. 2022), and pose disentanglement strategies (Tsai et al. 2022).
@@
Recently, inspired by the great achievement of BERT (Devlin et al. 2018) and MAE (He et al. 2021), many Transformer-based MPM methods for point cloud have been proposed and significantly outperform other methods in downstream tasks. Point-BERT (Yu et al. 2022) introduced the BERT pretraining scheme to point clouds, which trains models with a task to predict discrete tokens generated by dVAE (Rolfe 2016). Point-MAE (Pang et al. 2022) proposed a self-supervised pretext task to reconstruct masked input point patches based on the known tokens. Point-M2AE (Zhang et al. 2022) employed a pyramid encoder and decoder design to conduct masked point cloud reconstruction in a hierarchical manner. ACT (Dong et al. 2023), I2P-MAE (Zhang et al. 2023), ReCon (Qi et al. 2023) and ReCon++ (Qi et al. 2024) improved the 3D masked modeling by introducing pre-trained model from other modalities, such as images and natural languages. Point-GPT (Chen et al. 2024a) arranged point patches in an ordered sqeuence and extended the concept of genreative pre-training transformer (GPT) to point clouds. Despite achieving remarkable results, these MPM methods are sensitive to rotation, constraining their scalability and generalizability.
3 Method
3.1 Revisiting Masked Point Modeling
@@
We first revisit the masked point modeling (MPM) framework for point clouds. Existing methods, such as Point-MAE, segment the input point cloud $\mathbf{X}$ into $G$ point patches $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{G}$ using Farthest Point Sampling (FPS) and K-Nearest Neighborhood (KNN) algorithms. Subsequently, a masking ratio $\alpha$ is applied, partitioning the point patches into a visible set $\mathbf{P}_{v}=\{\mathbf{p}_{i}^{v}\}_{i=1}^{V}$ and a masked set $\mathbf{P}_{m}=\{\mathbf{p}_{i}^{m}\}_{i=1}^{M}$, where $V=(1-\alpha)\times G$ and $M=\alpha\times G$ are the numbers of visible and masked patches, respectively. Following this, a mini-PointNet (Qi et al. 2017a) is utilized to tokenize the visible point patches as point tokens $\mathbf{T}_{v}=\{\mathbf{t}_{i}^{v}\}_{i=1}^{V}$ and the coordinates of patch centers are projected as position embeddings $\{\mathbf{pos}_{i}\}_{i=1}^{G}$ via a learnable MLP (multi-layer perceptron). To this end, an encoding stage is introduced, during which only the visible tokens $\mathbf{T}_{v}$ are fed into a transformer-based encoder:
@@
3.2 Rotation-Invariant Transformer
@@
$\bar{\mathbf{p}}_{i},\mathbf{R}_{i}^{\prime}=\mathrm{PCA}(\mathbf{p}_{i}\mathbf{R}),\mathbf{R}_{i}^{\prime}=\mathbf{R}_{i}\mathbf{R},$ (7)
where $\mathbf{R}_{i}^{\prime}$ is the rotation that rotates the canonical $\bar{\mathbf{p}}_{i}$ into the rotated point patch $\mathbf{p}_{i}\mathbf{R}$. Hence, the orientation embedding $\mathbf{ori}_{i}$ is affected as well, since the MLP itself does not inherently ensure rotation invariance.
@@
where $\mathbf{Q}=\bar{\mathbf{T}}\mathbf{W}^{Q}$, $\mathbf{K}=\bar{\mathbf{T}}\mathbf{W}^{K}$, $\mathbf{V}=\bar{\mathbf{T}}\mathbf{W}^{V}$ are the query, key and value matrices; $\mathbf{W}^{Q},\mathbf{W}^{K}$ and $\mathbf{W}^{V}$ are learnable matrices; and $\mathbf{B}=\{b_{ij}\}\in\mathbb{R}^{G\times G}$ is RI-OE. We formulate $b_{ij}$ by establishing bilinear relations that integrate the features of the $i^{\text{th}}$ patch token with the relative orientation information pertaining to the $j^{\text{th}}$ patch:
@@
Note that for computational efficiency, we adopt the query weight matrix for the bilinear operator empirically. Instead of predicting the relative rotation with learnable module as done in existing method (Deng et al. 2023), our RI-OE directly employs rotation matrices to delineate the relative orientation, thereby avoiding intricate network designs.
@@
where $\mathbf{c}_{i}\mathbf{R}_{i}^{T}$ represents the location of the origin in the local reference frame of $\mathbf{p}_{i}$. The rotation invariance of $\mathbf{ripos}_{i}$ is discussed in the technical appendix.
@@
3.3 Rotation-Invariant MAE
Existing MPM methods utilize raw point clouds to guide the reconstruction of masked patches. However, this approach is not suitable for building a rotation-invariant MPM. Specifically, for a given point cloud in different directions, a rotation-invariant MPM should yield identical point cloud latent representations; however, the coordinates of the masked patches are different, necessitating that the decoder reconstructs distinct patches from the same latent representation. This conflict results in unstable loss, making the model difficult for optimization. We provide theoretical proof for this in the technical appendix.
@@ 3.3 Rotation-Invariant MAE
4 Experiments
4.1 Implementation Details
@@
For training details, we utilized an AdamW optimizer with a cosine learning rate decay, applying a decay factor 0.05, and incorporated a 10-epoch warm-up phase. The learning rate was set at 0.0005 for both pretraining and classification, while for segmentation, it was adjusted to 0.0002. We pretrained RI-MAE on ShapeNet (Chang et al. 2015), which consists of more than 50,000 3D models, covering 55 categories. Following ClusterNet (Chen et al. 2019), we conducted experiments in three scenarios to evaluate the rotation invariance: 1) z/z for training and testing with azimuthal rotations; 2) z/SO3 for training with azimuthal rotations and testing with arbitrary rotations; and 3) SO3/SO3, entailing both training and testing under arbitrary rotations. All experiments were conducted utilizing two RTX 2080Ti GPUs, with the PyTorch framework version 1.7.
4.2 Downstream Tasks
@@
We evaluate the proposed method on a classification task with a real-word dataset - ScanOjectNN (Uy et al. 2019), which contains 2,902 point clouds of 15 categories collected in real world. There are three variants of ScanOjectNN: OBJ_BG, OBJ_ONLY, and PB_T50_RS. We evaluate our approach on the OBJ_BG variant by incorporating additional MLP classification heads, with 1,024 points sampled, and compare our method with state-of-the-art rotation-invariant and MPM approaches. It is worth emphasizing that, for the MPM methods, experiments in rotation scenarios were conducted utilizing the open-source code. The results are detailed in Table 1. Specifically, we trained an RI-Transfomer without utilizing the proposed dual-branch self-supervised learning strategy as a baseline. The baseline RI-Transformer exhibits superior performance compared to the state-of-the-art rotation-invariant methods, notably outperforming the second-best approach (Yu, Zhang, and Cai 2023) by a margin of +1.7%. Regrading RI-MAE, with the dual-branch learning strategy, it delivers improvements of +2.6% over the RI-Transfomer baseline. These results indicate the superiority of the proposed mechanisms. The results of the remaining two variants are included in the technical appendix.
@@
ModelNet40 is a synthetic shape dataset that contains 12,311 CAD models across 40 categories, with 1,024 points sampled from each model (Wu et al. 2014). We conducted few-shot learning on ModelNet40, following practices from existing MPM methods. The performance is assessed across two settings: 10-way 10-shot and 10-way 20-shot, where $K$-way $N$-shot indicates that $N$ objects are sampled for training from each of $K$ randomly selected classes, followed by testing on 20 unseen instances randomly chosen from each of these $K$ classes. All experiments were conducted in the z/SO3 scenario to evaluate rotation robustness. We report both the average accuracy and the standard deviation from 10 independent runs in Table 2. The results show that our RI-MAE significantly outperforms existing methods, showing more than 50% improvements. We also compare performance in other train/test scenarios and report the results of 5-way 10-shot and 5-way 20-shot settings in the technical appendix.
@@
The evaluation on part segmentation is with ShapeNet, which consists of 16,881 objects across 16 categories. In line with existing MPM methods, we sample 2,048 points from each object as input. A simple segmentation head is incorporated as in Point-MAE (Pang et al. 2022). The segmentation results in the z/SO3 scenario are reported in Table 3. It is evident that existing MPM methods are sensitive to rotations, resulting in significantly unsatisfactory accuracy levels. In contrast, our RI-MAE demonstrates robustness to rotations and achieves the best performance, with a mean intersection over union (mIoU) score of 84.3% $\mathrm{mIoU}_{I}$ and 82.1% $\mathrm{mIoU}_{C}$. In addition, the part segmentation results of RI-MAE in the z/SO3 scenario are visualized in Fig. 3, indicating consistent predictions regardless of rotation. We also provide the results in the z/z scenario in the technical appendix.
@@
Table 5: Ablation study on RI-OE, RI-PE and dual-branch.